• A-Z
  • Directory
  • myUVM
  • Loading search...

Frank's Activity Log

14. Juni 2017: Mittwoch

Posted: June 15th, 2017 by fcs

Today’s Activities:

  • Backups:
    • DISK Clones completed yesterday at 18:20 (1d 3h 43m).
    • TAPE Clones completed yesterday at 20:19 (1d 5h 42m).
    • FSGen3 Clones completed at 7:12 (1d 16h 35m).
    • SN10 – Disk – 39R7348 arrived, installed, working.
  • LDAP/Accounts:
    • RPO sends a new person,
      Apparently, an accepted applicant who never attended
      and had a different name.
    • Processed the monthly retirees feed update.
    • Update the alumnus feed code – probably causing massive breakage.

13. Juni 2017: Dienstag

Posted: June 14th, 2017 by fcs

Today’s Activities:

  • Backups:
    • Stornode10 blows a disk at 1:51
      • 7979-71U 99P6286
      • IBM FRU 39R7348 (73.4GB 15K RPM SAS)
      • Park Place is sending a replacement.
    • Qualstar RMA # 50673:
      • New drive (1310146537, *502E2) arrived and installed.
      • First power on gave an SCD 5 – not good – second attempt worked.
      • Old drive sent back.
  • LDAP/Accounts:
    • PeopleSoft new employee missing… OH! no update last night.
      Extract file was not ready at the appointed hour.
    • HR says “Work Only” is the new privacy default – implement.
    • Updated account groups per request.
    • Start development of QD code to add uvmEduPurge objectClass to OLD entries.

12. Juni 2017: Montag

Posted: June 13th, 2017 by fcs

Today’s Activities:

  • Backups:
    • Checking on 130280L4 and 130245L4.
      • 130280L4(8LRVJA76) – 1133 GB, 9 savesets
      • 130245L4(8LRVKU0G) – 985 GB, 12 savesets
    • 9310145548, *526d7, B001T02E, SN9 ST0
      • Since both tapes have been vindicated,
        requesting yet another replacement drive from Qualstar.
      • Qualstar issued RMA # 50673
    • Reclaimed two (2) tapes for the vaults.
    • Added forty (40) tapes to the Scratch pool.
    • BUJBOD1, 32:3 (ST4000NM0023, A001, Z1Z0AP1M) failed.
      • Replaced with ST4000NM0025, N002, ZC10FFXX
    • Clones started.
    • Maintenance announced for June 19.
  • LDAP/Accounts:
    • Resources process goes KABOOM – There was an earth shattering KA-BOOM!
      • Updated the perl script to ignore SIP and X500 type (all non-SMTP type)
        entries in the proxyAddresses returned from AD.
    • Oh look – The same non-SMTP values are in the PEOPLE entries too.
      • Updated the perl script to use the Add= and Remove=
        subcommands on set-Mailbox’s -emailAddresses argument.

9. Juni 2017: Freitag

Posted: June 12th, 2017 by fcs

Today’s Activities:

  • Backups:
    • The tape drive in E02 throws SCD 6. /sigh
      • SN9 ST0, 9310145548, *526d7, B001T02E
      • ITDT dump taken, sent to Qualstar for analysis.
      • Drive disabled in NetWorker.
      • Qualstar says they think it is the tape…
        • Timeline
        • Drive powered on 6/6 ~11:00
        • First failure: 120539L4 @ 6/6 11:53 – premature EOT.
        • Second failure: 120621L4 @ 6/7 17:41 – premature EOT.
        • Third failure: 130280L4 @ 6/8 6:29 – Tape Alert – Write failure.
        • Fourth failure: 130245L4 @ 6/9 6:57 – Tape Alert – Read failure.
        • Drive appears to stop communicating for a while – several timeout errors.
      • 8LRVJA76 = 130280L4; 8LRVKU0G = 130245L4
      • Stage data off these two tapes, let them be reused.
      • Leave drive disabled for the weekend.
      • If the tapes have no issues over the weekend, get new drive.
      • Staging of data from 130280L4 caused 21 consecutive read errors.
        Not surprising since there was a “TAPE ALERT” “WRITE ERROR”
      • Labeled both tapes, so they are ready to be used.
    • FSGen3 clones finished on June 7 at 15:23 (2d 0h 33m).
    • Tape Run:
      • Forty (40) tapes taken.
      • Forty (40) tapes returned.
      • Forty (40) tapes labeled.
    • Define three (3) physical clients for EAS.
      But they were on a network that is not allowed to talk to
      ozzie, so I had to punch more holes in the firewall.
  • LDAP/Accounts:
    • OpsGenie account was restored in AD to where it belongs.

8. Juni 2017: Donnerstag

Posted: June 9th, 2017 by fcs

Today’s Activities:

  • Backups:
    • stornode2’s staging process borked on Tuesday – clean up the mess.
  • LDAP/Accounts:
    • LDAP->AD update borked. Clean up the mess.

7. Juni 2017: Mittwoch

Posted: June 9th, 2017 by fcs

Today’s Activities:

  • Backups:
    • DISK clones completed at 01:15 (1d 10h 26m).
    • Added forty (40) tapes to the Scratch pool.
  • LDAP/Accounts:
    • Fixed BUG 20170606.
    • Cleanup stub ldap entry from 3+ years ago.
    • Repair cisco ACLs – too restrictive.

6. Juni 2017: Dienstag

Posted: June 7th, 2017 by fcs

Today’s Activities:

  • Backups:
    • Qualstar RMA # 50660 (131009051, *450ab, E02, SN9 ST0:
      • New tape drive arrived.
      • New tape drive installed.
      • 9310145548, *526d7
    • BUJBOD1 54:5, 32:3, 54:3:
      • Box of disks arrived.
      • 54:5 Z1Z08FZ8 replaced with ZC11PWYC.
      • Waiting for copyback to complete before replacing others.
    • COMIS clones completed yesterday at 17:46 (2h 57m).
  • LDAP/Accounts:
    • cisco VOIP ACL changes pushed out to production.
    • Testing update code.
    • Bad code update… caused failure.
    • A stray $ in data from Banner uncovered a lack of defense.
  • Footprints:
    • Several of the accounts actually expired in AD.
    • Had to have them marked to never expire.

5. Juni 2017: Montag

Posted: June 6th, 2017 by fcs

Today’s Activities:

  • Backups:
    • BUJBOD1 Disk 54:5 failed on Saturday:
      • Sonntag: ran lsiget, sent the output to AMS Storage.
    • Realize that I read the report wrong and the two drives with SMART
      IMPENDING FAILURE warnings in BUJBOD1 are 32:3 and 54:3.
    • AMS Storage is sending three drives. YAY!
    • Reclaim two (2) tapes for the vaults.
    • The tape drive *450ab installed on Friday has thrown errors all weekend.
      • 131009051, *450ab, E02, SN9 ST0 – Multiple SCD 6’s
      • Disabled the drive.
      • ITDT taken, sent to Qualstar.
      • RMA # 50660 issued – this drive was DOA.
  • LDAP/Accounts:
    • Registrar’s office approves proposal to clean up former student
      entries that have no backing for three months.
    • Converted the HL accounts to be departmental instead of miscellaneous type
    • CUCM requires more changes ACL and individual entries.
      • Testing changes… will roll in production later.
    • Testing update code that does not remove FS kerb principal for
      manual lifetime-email entitlement.

2. Juni 2017: Freitag

Posted: June 5th, 2017 by fcs

Today’s Activities:

  • Backups:
    • SB group still crawling from last weekend – sad.
    • Tape Vaulting:
      • 61 tapes need to go to vaults.
      • 62 tapes ejected and taken to vaults.
      • 80 tapes returned and labeled.
    • Split the crawling SB group in two so the slow system can mess up just itself.
    • Re-writing the tobeheld script so it puts tapes into Scratch in the
      correct order.
    • DISK clones completed at 11:01 (3d 3h 11m).
  • LDAP/Accounts:
    • OpenLDAP 2.4.45 is released upon the world…
      • I need to download, build, test.
      • Downloaded, built, testing.
    • Got a new RPO person with a bad surname – found, set up to resolve tonight.
    • Responding to RPO about the failure from last night.
    • Tweaking accounts for use testing voip phones

1. Juni 2017: Donnerstag

Posted: June 2nd, 2017 by fcs

Today’s Activities:

  • Backups:
    • Qualstar RMA # 50654: *5b0cb, 1310186868, SN4 ST0:
      • Drive replaced.
      • Old drive delivered to mail room for shipping.
    • One of the clients removed yesterday was actually needed!
      • goat needs netfiles01 to be defined – why???
      • nsrclientfix renamed the “netfiles01” client to be “goat” causing
        duplication, confusion and failure.
      • Again, I am shown that EMC’s advice is not to be taken at face value.
    • Added forty (40) tapes to the Scratch pool.
    • 1310234952, *54E20, SN9 ST0: SCD A
      • This drive was installed on May 24… 7 days!
      • ITDT dump taken and sent to Qualstar for examination.
      • Qualstar issues RMA # 50655 to send this one back.
    • Reached out to AMS Storage about:
      • Renewal pricing for support of JBOD with expiring warranty.
      • Time of extended support on the JBOD we renewed last year.
      • Status on the two LSIGET’s I sent them in the past week.
    • FSGen3 clones finished at 7:30 (1d 23h 39m).
    • Purged the three invalid duplicate clone save sets (WiSS Conversion issue).
    • More work on the backup rfi …
  • LDAP/Accounts:
    • Former Student realm audit finds problem – fixed.
    • Developing proposal for clearing out place holder entries in branches.
Contact Us ©2010 The University of Vermont – Burlington, VT 05405 – (802) 656-3131
Skip to toolbar